OpenStudy (anonymous):

Solving Linear system with three variables 2x-3y+z=4 5x+2y+2z=2 x+4y-3z=7 how do i start

OpenStudy (anonymous):

ok , tell me what letter easy to cancel ? "z" right ?

OpenStudy (anonymous):

we have 2x-3y + z = 4 ( only z in the function ) so we move all them to the right side , except z ....can you do it ? show me how to do, i will show you next step

OpenStudy (anonymous):

2x-3y+z=4 *2 2x*2-3y*2+2*z=2*4 = 4x-6y+2z=8 add 5x+2y+2z=2 = 9x-4y+16z=10